[figure: 9.7 cm]

INDUCED BY MAGNETIC FIELD PHASE TRANSITIONS IN

NEODYMIUM MANGANETES MONOCRYSTALS

Yu.F. Popov^{a}, A.M. Kadomtseva^{a}, K.I. Kamilov^{a}, G.P. Vorob’ev^{a}, J.S.

Shtofich^{a}, A.A. Mukhin^{b}, V.Yu. Ivanov^{b}, A.M. Balbashov^{c} aMoscow State University, 119992 Moscow, Russian Federation

bGeneral Physics Institute of the Russian Acad. Sci., 38 Vavilov St., 119991 Moscow,

Russia, Moscow Power Engineering Institute, 105835 Moscow, Russia

Magnetic, anisotropic and magnetoelastic properties have been investigated for NdMnO_{3} and Nd_{0.5}Sr_{0.5}MnO_{3} monocrystals. in pulsed magnetic fields up to 250 kOe in temperature range 10-300 K. Phase transitions induced by strong magnetic fields H_{C}= 100 kOe have been discovered. These transitions have been accompanied by sharp anomalies of magnetization and magnetostriction and linked with reorientation of magnetic moment from c to b axis of rhombic crystal. Critical fields that induce spin reorientation decreased with temperature decreasing because of magnetization increasing accords to Nd-Mn exchange. Corresponding H-T phase diagrams were obtained. Also induced by magnetic field phase transitions in monocrystal Nd_{0.5}Sr_{0.5}MnO_{3} have been investigated. Nd_{0.5}Sr_{0.5}MnO_{3} have ferromagnetic metallic state under T_{C}=250 K at T¡T_{CO}=140 K antiferromagnetic charge ordering state. For Nd_{0.5}Sr_{0.5}MnO_{3} monocrystals jumps, linked with suppression of charge ordering, of magnetization and magnetostriction have been observed in temperature range 10 -140K. Phase H-T diagrams have been obtained. This work is support by RFBR (00-15-96695 and 00-02-16500).

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Yu.F. Popov

Address for correspondence : Department of Physics Moscow State University Leninskie Gori, 119992 Moscow, Russian Federation

Email address : popov@plms.phys.msu.su